# Boundary-driven quantum systems near the Zeno limit: steady states and long-time behavior

### v2

@@ 1 Introduction
$\|e^{t\mathcal{L}_{\gamma}}\rho_{0}-\pi_{A}\otimes e^{t\mathcal{L}_{P,\gamma}}R_{0}\|_{1}$ $\leqslant$ $\|e^{t\mathcal{L}_{\gamma}}\rho_{0}-e^{t\mathcal{L}_{\gamma}}\pi_{A}\otimes R_{0}\|_{1}$
$+$ $\|e^{t\mathcal{L}_{\gamma}}\pi_{A}\otimes R_{0}-\pi_{A}\otimes R(t)\|_{1}$
$+$ $\|\pi_{A}\otimes R(t)-\pi_{A}\otimes e^{t\mathcal{L}_{P,\gamma}}R_{0}\|_{1}\ .$
@@
The situation near, but not at, the Zeno limit is more interesting, and then there is dissipation on time scales of order $\gamma$, and the more precise bound (16) can be used to realte the rates of approach to stationarity for $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$. However, the analysis of both of these Lindblad generators is complicated because in both the coherent and dissipative effects operate on different times scales. For instance, with $\mathcal{L}_{P,\gamma}$ the coherent evolution due to $\mathcal{K}_{P}$ acts on time scales of order one, while the dissipative evolution due to $\mathcal{D}_{P}$ acts on time scale $\mathcal{O}(\gamma)$.
@@
The results (22) and (23) allow us to study the evolutions governed by $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ in terms of the evolution governed by $\mathcal{D}_{P}^{\sharp}$. In particular, we use (22) and (23) to show that is $\mathcal{D}_{P}^{\sharp}$ is egodic and gapped, then so are $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ for all sufficiently large $\gamma$, and moreover, we prove a tight relation among the rates of aproach to stationairty for all three generators.
This is best done in terms of the mixing time, which plays an important role in several recent investigations of open quantum systems [15, 33, 34, 35]. We recall the relevant definitions:
@@
with the convention that $t_{{\rm mix}}(\mathcal{L},\epsilon)=\infty$ if there is no finite time $t$ such that $d_{{\rm TV}}(e^{t\mathcal{L}}\rho_{0},e^{t\mathcal{L}}\rho_{1})<\epsilon$ for all pairs density matrices $\rho_{0},\rho_{1}$. It is known that if $t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$ for some $0<\epsilon<\tfrac{1}{2}$, then this is true for all such $\epsilon$; see Theorem E.3.
@@
Returning to $\mathcal{L}_{\gamma}$, $\mathcal{L}_{P,\gamma}$ and $\mathcal{D}_{P}^{\sharp}$, we prove in Theorem 7.5 that for all but at most contably many values of $\epsilon_{0}\in(0,\tfrac{1}{2})$,
@@
This shows one way in which $\mathcal{D}_{P}^{\sharp}$ governs the long-time behavior of the evolution described by both $\mathcal{L}_{P,\gamma}$ , and by $\mathcal{L}_{\gamma}$ , near the Zeno limit. For another, suppose that $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped. and let $\bar{R}$ be the unique density matrix on $\mathcal{H}_{B}$ such that $\mathcal{D}_{P}^{\sharp}\bar{R}=0$. Let $\bar{\rho}_{\gamma}$ be the unique (for $\gamma$ sufficiently large) density matrix on $\mathcal{H}_{AB}$ satisfying $\mathcal{L}_{\gamma}\bar{\rho}_{\gamma}=0$. Then we prove that
@@
There is a large literature on the perturbative approach; see [20, Section III.D]. The simplest cases concern a Linblad equation of the type
@@
in which $\mathcal{L}_{0}$ is ergodic so that there is a unique denisty matrix $\bar{\rho}_{0}$ satisfying $\mathcal{L}_{0}\bar{\rho}_{0}=0$. One then assumes that for $\epsilon$ sufficiently small, there is a unique density matrix $\bar{\rho}_{\epsilon}$ satisfying $(\mathcal{L}_{0}+\epsilon\mathcal{L}_{1})\bar{\rho}_{\epsilon}=0$ and moreover (the key assumption), that $\bar{\rho}_{\epsilon}$ has an expansion
@@
Our problem is more complex. Here, $\mathcal{D}$ plays the role of $\mathcal{L}_{0}$ and it is far from ergodic. Therefore, there is no obvious candidate for $\bar{\rho}_{0}$. A related problem arises in [21] for a boundary driven spin chain model (but not near the Zeno limit). In that case, a conjectured ansatz [21, eq. (6)] based on exact calculation for short chains provides starting point $\rho_{0}$, but there is no proof that is the correct starting point of the expansion.
@@ 2 Notation and some mathematical preliminaries
for all $Z$ in $\widehat{\mathcal{H}}$. In our analysis, this plays the role of the “vectorization” discussed in [20, Section III.B], but avoids identifying $\widehat{\mathcal{H}}$ with $\mathcal{H}\otimes\mathcal{H}$ and $\widehat{\mathcal{K}}$ with $\mathcal{K}\otimes\mathcal{K}$ which requires a choice of basis.
@@ 4 The super-operators $\mathcal{K}_{P}$ and $\mathcal{D}_{P}$
The inequality (61), derived in this manner, also appears in Lindblad’s work [23, (3.1)] on the generators of CPTP semigroups. Now apply (61) as follows:
@@ 8 Application to the computation of steady states
Under the assumption the $\mathcal{D}_{P}^{\sharp}$ is ergodic, there are is a uniquely determined sequence $\{\bar{n}_{k}\}_{k\geqslant 0}$ in $\widehat{\mathcal{H}}_{B}$ such that the equations (124) and (125) are all satisfied, and moreover for all $k\geqslant 0$,
$\|\bar{n}_{k}\|_{1}\leqslant 2(\|\mathcal{S}\mathcal{K}\|_{1\to 1}+(C_{\mathcal{V}}+C_{\mathcal{W}}))^{k}\left(\|\mathcal{S}\mathcal{K}\|_{1\to 1}+\frac{\|\mathcal{D}_{P}\bar{R}\|_{1}}{b}\right)\ .$ (136)
@@
Consequently for $\gamma>(\|\mathcal{S}\mathcal{K}\|_{1\to 1}+(C_{\mathcal{V}}+C_{\mathcal{W}}))$, $\sum_{k=0}^{\infty}\gamma^{-k}n_{k}$} converges in the trace norm, and hence
@@ 9 Hydrodynamic limits
A generalization of the super-operator trace norm plays an important role in quantum information theory because of its connections with quantum entropy. For $1\leqslant p<\infty$ the Schatten $p$-norm on $\widehat{\mathcal{H}}$ is defined by $\|X\|_{p}=\left(\textnormal{Tr}[(X^{*}X)^{p/2}]\right)^{1/p}$. For $p=2$ this is simply the Hilbert space norm associated to the Hilbert-Schmidt inner product. For $p=1$, this is the trace norm. For all $X$, $\|X\|_{p}$ is monotone non-increasing, and $\lim_{p\to\infty}\|X\|_{p}=\|X\|_{\infty}$ where $\|X\|_{\infty}$ is the operator norm of $X$, which is equal to its largest singular value of $X$.
Let $\mathcal{H}$ be a finite dimensional Hilbert space, and let $\mathcal{T}$ be a linear operator on $\widehat{\mathcal{H}}$ For $1\leqslant p\leqslant\infty$. Recall the definition (33) of the super-operator trace norm:
$\|\mathcal{T}\|_{1\to p}=\sup\{\|\mathcal{T}(X)\|_{p}\ :\|X\|_{1}=1\ \}\ .$
We need a result of a similar nature. Let $\mathcal{D}_{A}$ be a Lindblad generator on $\widehat{\mathcal{H}}_{A}$, and let $\mathcal{D}:=\mathcal{D}_{A}\otimes\mathds{1}_{B}$, as in the introduction. Let $\mathcal{Q}_{A}$ and $\mathcal{Q}$ be the projections onto the ranges of $\mathcal{D}_{A}$ and $\mathcal{D}$ respectively defined in (37) and (40). We need to know that when $\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\|_{1\to 1}$ is small, then so is
$\|e^{t\mathcal{D}}\mathcal{Q}_{A}\|_{1\to 1}=\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\otimes\mathds{1}_{B}\|_{1\to 1}\ .$
While $e^{t\mathcal{D}_{A}}$ is CPTP, $e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}$ is not; it is instead the diference between two CPTP maps, $e^{t\mathcal{D}_{A}}$ and $e^{t\mathcal{D}_{A}}\mathcal{P}_{A}$. Thus we cannot apply the bound (152) to conclude that $\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}=\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\|_{1\to 1}$.
The following lemma will however allow us to transfer a certain bound on $\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\|_{1\to 1}$ directly to $\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}$, thus bounding the latter that is independent of the dimension of $\mathcal{H}_{B}$.

### v3

@@ 1 Introduction
$\|e^{t\mathcal{L}_{\gamma}}\rho_{0}-\pi_{A}\otimes e^{t\mathcal{L}_{P,\gamma}}R_{0}\|_{1}$ $\leqslant$ $\|e^{t\mathcal{L}_{\gamma}}\rho_{0}-e^{t\mathcal{L}_{\gamma}}\mathcal{P}\rho_{0}\|_{1}$
$+$ $\|e^{t\mathcal{L}_{\gamma}}\pi_{A}\otimes R_{0}-\mathcal{P}(e^{t\mathcal{L}_{\gamma}}\pi_{A}\otimes R_{0})\|_{1}$
$+$ $\|\mathcal{P}(e^{t\mathcal{L}_{\gamma}}\pi_{A}\otimes R_{0})-\pi_{A}\otimes e^{t\mathcal{L}_{P,\gamma}}R_{0}\|_{1}\ .$
@@
The situation near, but not at, the Zeno limit is more interesting, and then there is dissipation on time scales of order $\gamma$, and the more precise bound (16) can be used to relate the rates of approach to stationarity for $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$. However, the analysis of both of these Lindblad generators is complicated because in both the coherent and dissipative effects operate on different times scales. For instance, with $\mathcal{L}_{P,\gamma}$ the coherent evolution due to $\mathcal{K}_{P}$ acts on time scales of order one, while the dissipative evolution due to $\mathcal{D}_{P}$ acts on time scale $\mathcal{O}(\gamma)$.
@@
The results (22) and (23) allow us to study the evolutions governed by $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ in terms of the evolution governed by $\mathcal{D}_{P}^{\sharp}$. In particular, we use (22) and (23) to show that if $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped, then so are $\mathcal{L}_{\gamma}$ and $\mathcal{L}_{P,\gamma}$ for all sufficiently large $\gamma$, and moreover, we prove a tight relation among the rates of approach to stationarity for all three generators.
This is best done in terms of the mixing time which plays an important role in several recent investigations of open quantum systems [15, 35, 36, 37]. We recall the relevant definitions:
@@
with the convention that $t_{{\rm mix}}(\mathcal{L},\epsilon)=\infty$ if there is no finite time $t$ such that $d_{{\rm TV}}(e^{t\mathcal{L}}\rho_{0},e^{t\mathcal{L}}\rho_{1})<\epsilon$ for all pairs of density matrices $\rho_{0},\rho_{1}$. It is known that if $t_{{\rm mix}}(\mathcal{L},\epsilon)<\infty$ for some $0<\epsilon<\tfrac{1}{2}$, then this is true for all such $\epsilon$; see Theorem E.3.
@@
Returning to $\mathcal{L}_{\gamma}$, $\mathcal{L}_{P,\gamma}$ and $\mathcal{D}_{P}^{\sharp}$, we prove in Theorem 7.5 that for all but at most countably many values of $\epsilon_{0}\in(0,\tfrac{1}{2})$,
@@
This shows one way in which $\mathcal{D}_{P}^{\sharp}$ governs the long-time behavior of the evolution described by both $\mathcal{L}_{P,\gamma}$ , and by $\mathcal{L}_{\gamma}$ , near the Zeno limit. For another, suppose that $\mathcal{D}_{P}^{\sharp}$ is ergodic and gapped, and let $\bar{R}$ be the unique density matrix on $\mathcal{H}_{B}$ such that $\mathcal{D}_{P}^{\sharp}\bar{R}=0$. Let $\bar{\rho}_{\gamma}$ be the unique (for $\gamma$ sufficiently large) density matrix on $\mathcal{H}_{AB}$ satisfying $\mathcal{L}_{\gamma}\bar{\rho}_{\gamma}=0$. Then we prove that
@@
There is a large literature on the perturbative approach; see [22, Section III.D]. The simplest cases concern a Lindblad equation of the type
@@
in which $\mathcal{L}_{0}$ is ergodic so that there is a unique density matrix $\bar{\rho}_{0}$ satisfying $\mathcal{L}_{0}\bar{\rho}_{0}=0$. One then assumes that for $\epsilon$ sufficiently small, there is a unique density matrix $\bar{\rho}_{\epsilon}$ satisfying $(\mathcal{L}_{0}+\epsilon\mathcal{L}_{1})\bar{\rho}_{\epsilon}=0$ and moreover (the key assumption), that $\bar{\rho}_{\epsilon}$ has an expansion
@@
Our problem is more complex. Here, $\mathcal{D}$ plays the role of $\mathcal{L}_{0}$ and it is far from ergodic. Therefore, there is no obvious candidate for $\bar{\rho}_{0}$. A related problem arises in [23] for a boundary driven spin chain model (but not near the Zeno limit). In that case, a conjectured ansatz [23, eq. (6)] based on exact calculation for short chains provides starting point $\rho_{0}$, but there is no proof that is the correct starting point of the expansion.
@@ 2 Notation and some mathematical preliminaries
for all $Z$ in $\widehat{\mathcal{H}}$. In our analysis, this plays the role of the “vectorization” discussed in [22, Section III.B], but avoids identifying $\widehat{\mathcal{H}}$ with $\mathcal{H}\otimes\mathcal{H}$ and $\widehat{\mathcal{K}}$ with $\mathcal{K}\otimes\mathcal{K}$ which requires a choice of basis.
@@ 4 The super-operators $\mathcal{K}_{P}$ and $\mathcal{D}_{P}$
The inequality (61), derived in this manner, also appears in Lindblad’s work [25, (3.1)] on the generators of CPTP semigroups. Now apply (61) as follows:
@@ 8 Application to the computation of steady states
Under the assumption the $\mathcal{D}_{P}^{\sharp}$ is ergodic, there is a uniquely determined sequence $\{\bar{n}_{k}\}_{k\geqslant 0}$ in $\widehat{\mathcal{H}}_{B}$ such that the equations (124) and (125) are all satisfied, and moreover for all $k\geqslant 0$,
$\|\bar{n}_{k}\|_{1}\leqslant 2(\|\mathcal{S}\mathcal{K}\|_{1\to 1}+(C_{\mathcal{V}}+C_{\mathcal{W}})^{k}\left(\|\mathcal{S}\mathcal{K}\|_{1\to 1}+\frac{\|\mathcal{D}_{P}\bar{R}\|_{1}}{b}\right)\ .$ (136)
@@
Consequently for $\gamma>(\|\mathcal{S}\mathcal{K}\|_{1\to 1}+(C_{\mathcal{V}}+C_{\mathcal{W}})$, $\sum_{k=0}^{\infty}\gamma^{-k}n_{k}$} converges in the trace norm, and hence
@@ 9 Hydrodynamic limits
In Appendix C we shall use the spectral gap of $\mathcal{D}_{A}$ to bound the rate of decay of $\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\|_{1\to 1}=\|e^{t\mathcal{D}_{A}}-|\pi_{A}\rangle\langle\mathds{1}_{A}|\|_{1\to 1}$. However, we need control on the rate at which $\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}=\|e^{t\mathcal{D}}-\mathcal{P}\|_{1\to 1}$ tends to zero. Since $e^{t\mathcal{D}}=e^{t\mathcal{D}_{A}}\otimes\mathds{1}_{B}$ and $\mathcal{P}=|\pi_{A}\rangle\langle\mathds{1}_{A}|\otimes\mathds{1}_{B}$, this might seem automatic, but it is not. If $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are two CPTP super-operators on $\mathcal{H}_{A}$, then it can be the case that
$\|(\mathcal{T}_{1}-\mathcal{T}_{2})\otimes\mathds{1}_{B}\|_{1\to 1}>\|\mathcal{T}_{1}-\mathcal{T}_{2}\|_{1\to 1}\ .$
This has been investigated by Kitaev [20, 21] who introduced the diamond norm on Hermitian super-operators; ie.e, those for which $\mathcal{T}(X^{*})=\mathcal{T}(X)^{*}$ for all $X$. By a well-known variant of Choi’s Theorem, every Hermitian super-operator $\mathcal{T}$ can be written as the difference of two CPTP maps; see, e.g. [38]. Kitaev [20] proved that while it can be the case that $\|(\mathcal{T}_{1}-\mathcal{T}_{2})\otimes\mathds{1}_{B}\|_{1\to 1}$ increases with the dimension of $\mathcal{H}_{B}$, there is no further increase after the dimension of $\mathcal{H}_{B}$ has reached the dimension of $\mathcal{H}_{A}$. The diamond norm of $\mathcal{T}_{1}-\mathcal{T}_{2}$ is then defined by
$\|\mathcal{T}_{1}-\mathcal{T}_{2}\|_{\diamond}:=\|(\mathcal{T}_{1}-\mathcal{T}_{2})\otimes\mathds{1}_{B}\|_{1\to 1}\quad{\rm where}\quad{\rm dim}(\mathcal{H}_{B})={\rm dim}(\mathcal{H}_{A})\ .$
However, this leaves open the question of how much larger $\|\mathcal{T}_{1}-\mathcal{T}_{2}\|_{\diamond}$ might be than $\|\mathcal{T}_{1}-\mathcal{T}_{2}\|_{1\to 1}$. In our finite dimensional setting, all norms are equivalent, so hence if $\lim_{t\to\infty}\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}=0$, then $\lim_{t\to\infty}\|e^{t\mathcal{D}}\mathcal{Q}\|_{\diamond}=0$, and evidently $\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}\leqslant\|e^{t\mathcal{D}}\mathcal{Q}\|_{\diamond}$. However, there do not seem to be good general bounds that can be used to compare the rates.
The following lemma will however allow us to transfer a certain bound on $\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\|_{1\to 1}$ directly to $\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}$, and this bound of the latter is independent of the dimension of $\mathcal{H}_{B}$.
@@
In other words, for rank-one super-operators $\||Y\rangle\langle X|$, $\||Y\rangle\langle X|\|_{\diamond}=\||Y\rangle\langle X|\|_{1\to 1}$.
@@
There is a close connection between the issue discussed in this appendix with the norm multiplicitivity problem, and we conclude by briefly discussing the connection. A generalization of the super-operator trace norm plays an important role in quantum information theory because of its connections with quantum entropy. For $1\leqslant p<\infty$ the Schatten $p$-norm on $\widehat{\mathcal{H}}$ is defined by $\|X\|_{p}=\left(\textnormal{Tr}[(X^{*}X)^{p/2}]\right)^{1/p}$. For $p=2$ this is simply the Hilbert space norm associated to the Hilbert-Schmidt inner product. For $p=1$, this is the trace norm. For all $X$, $\|X\|_{p}$ is monotone non-increasing, and $\lim_{p\to\infty}\|X\|_{p}=\|X\|_{\infty}$ where $\|X\|_{\infty}$ is the operator norm of $X$, which is equal to its largest singular value of $X$.
Let $\mathcal{H}$ be a finite dimensional Hilbert space, and let $\mathcal{T}$ be a linear operator on $\widehat{\mathcal{H}}$. For $1\leqslant p\leqslant\infty$, define
$\|\mathcal{T}\|_{1\to p}=\sup\{\|\mathcal{T}(X)\|_{p}\ :\|X\|_{1}=1\ \}\ .$
While $e^{t\mathcal{D}_{A}}$ is CPTP, $e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}$ is not; it is instead the diference between two CPTP maps, $e^{t\mathcal{D}_{A}}$ and $\mathcal{P}_{A}$. Thus we cannot apply the bound (152) to conclude that $\|e^{t\mathcal{D}}\mathcal{Q}\|_{1\to 1}=\|e^{t\mathcal{D}_{A}}\mathcal{Q}_{A}\|_{1\to 1}$, and based on Kitaev’s results, such an identity is unlikely to be true in general. However, Lemma A.1 will give us what we need in Appendix C.